OpenStudy (ashleyvaldes501):

What is the equation in point-slope form of the line passing through (−2, 0) and (2, 8)? y = 2(x + 2) y = 2(x − 2) y = −2(x − 2) y = −2(x + 2)

OpenStudy (calculusxy):

Point-slope form: \(\large y - y_1 = m(x - x_1)\) Find the slope: \(\large \frac{y_2 - y_1}{x_2 - x_1}\) Have \(y_1 = -2 \) and \(x_1 = 0\). The \(y\) and \(x\) in the formula will just stay as it is.

OpenStudy (jdoe0001):

\(\begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -2}}\quad ,&{\color{blue}{ 0}})\quad % (c,d) &({\color{red}{ 2}}\quad ,&{\color{blue}{ 8}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (sshayer):

\[x _{1}=-2,y _{1}=0,x _{2}=2,y _{2}=8\] eq. of a line through two points is \[y-y_{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\left( x-x _{1} \right)\]

OpenStudy (sshayer):

you can take any point as (x1,y1) and other as (x2,y2)